OpenStudy (anonymous):

Evaluate[\ \lim_{n \to \infty }\, \frac{(2n-1)!}{(2n+1)!}\]

OpenStudy (anonymous):

\[\lim_{n \to \infty }\, \frac{(2n-1)!}{(2n+1)!}\]

myininaya (myininaya):

Try rewriting the bottom factorial

myininaya (myininaya):

\[(2n+1)!=(2n+1)(2n) \cdot (2n-1)!\]

myininaya (myininaya):

Let me know if you still need help on this one.

OpenStudy (anonymous):

Thanks myin

myininaya (myininaya):

Np. I hope you see something cancels. What do you think the answer is?

OpenStudy (anonymous):

I got 0

myininaya (myininaya):

0 is right since the degree on bottom is bigger than the degree on top